You’ve observed the following returns on SkyNet Data Corporation’s stock over the past five years: 20 percent, –12 percent, 17 percent, 20 percent, and 10 percent. a. What was the arithmetic average return on the company's stock over this five-year period? (Do not round intermediate calculations and enter your answer as a percent rounded to 1 decimal place, e.g., 32.1.) b-1. What was the variance of the company's returns over this period? (Do not round intermediate calculations and round your answer to 5 decimal places, e.g., .16161.) b-2. What was the standard deviation of the company’s returns over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a) Arithmetic Return =(20%-12%+17%+20%+10%)/5 =11%

b-1) Variance =((20%-11%)^2+(-12%-11%)^2+(17%-11%)^2+(20%-11%)^2+(10%-11%)^2)/(5-1) =1.82%

b-2) Standard Deviation =Variance ^0.5 =1.82%^0.5 =13.49%